OpenStudy (anonymous):

Solve for x: negative 7 over 6, multiplied by x minus 6 equals negative 48

OpenStudy (anonymous):

Hello, is your question -7/6 (x-6) = -48? If so, you first have to isolate x (get x by itself) and then solve the equation normally. to get rid of the -7/6 you simply multiply both sides of the equation by -6/7 which cancels it out. here you go -7/6(x-6) = -48 -6/7* -7/6 (x-6) = -48 * (-6/7) x-6 = 288/7 +6 +6 x=288/7 + 42/7 x=330/7 hope that helps!